Billy likes to go cycling.
His bike has wheels of diameter 85cm.
He has invented a counter for his bike,
which counts the number of revelations the wheel makes.
One day the counter shows 240 revolutions.
How far has billy cycled?
give your answer to the nearest metre.

Answer by Alan3354(69443)

His bike has wheels of diameter 85cm.
He has invented a counter for his bike,
which counts the number of revelations the wheel makes.
One day the counter shows 240 revolutions.
How far has billy cycled?
give your answer to the nearest metre.
---------------
Circumference = pi*85 cm
d = 240*85*pi = 20400*pi cm = 204*pi meters
d =~ 641 meters
------------------
You can buy those for ~ $20, they show speed and distance. No need to invent it. It works well on bicycles, but adapting it to rollerblades is not easy.